OpenStudy (anonymous):

hi how to prove e^i(pi)=-1

OpenStudy (kc_kennylau):

\[\Large\begin{array}{llr} &e^{i\pi}\\ =&\cos\pi+i\sin\pi&\mbox{(Euler's formula)}\\ =&-1+0i\\ =&-1 \end{array}\]

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Euler's_formula

OpenStudy (kainui):

That kind of depends on if you believe euler's magical looking formula is true or not. I barely believe it myself, why should there be any relation between euler's number and the quadratic forms? But if you expand out e^x, sinx, and cosx into infinite polynomials it makes it more obvious when you plug in i. Weird stuff.

OpenStudy (anonymous):

\[\begin{align*}e^{ix}&=\sum_{n=0}^\infty \frac{(ix)^n}{n!}\\\\\\ &=1+ix+\frac{i^2x^2}{2!}+\frac{i^3x^3}{3!}+\frac{i^4x^4}{4!}+\frac{i^5x^5}{5!}+\frac{i^6x^6}{6!}+\cdots\\ &=\left(1+\frac{i^2x^2}{2!}+\frac{i^4x^4}{4!}+\frac{i^6x^6}{6!}+\cdots\right)+\left(ix+\frac{i^3x^3}{3!}+\frac{i^5x^5}{5!}+\cdots\right)\\ &=\left(1+\frac{i^2x^2}{2!}+\frac{i^4x^4}{4!}+\frac{i^6x^6}{6!}+\cdots\right)+i\left(x+\frac{i^2x^3}{3!}+\frac{i^4x^5}{5!}+\cdots\right)\\ &=\left(1-\frac{x^2}{2!}+\frac{x^4}{4!}-\frac{x^6}{6!}+\cdots\right)+i\left(x-\frac{x^3}{3!}+\frac{x^5}{5!}-\cdots\right)\\\\\\ &=\sum_{n=0}^\infty \frac{(-1)^{n}x^{2n}}{(2n)!}+i\sum_{n=0}^\infty \frac{(-1)^nx^{2n+1}}{(2n+1)!}\\\\\\ &=\cos x+i\sin x \end{align*}\] So, when \(x=\pi\), you get \(e^{i\pi}=\cos \pi+i\sin \pi=-1+i\times 0=-1\).

OpenStudy (kainui):

A fun little extra: \[e^x=e^{i(-ix)}=\cos(ix)-isin(ix)=\cosh(x)+\sinh(x)\]

OpenStudy (kc_kennylau):

@Kainui Wow I didn't know this thanks :D

OpenStudy (anonymous):

never thought of mclaurin's thanks a loads, @SithangGiggles